OpenStudy (anonymous):

Solve 7x - 2y = -3 14x + y = 14

OpenStudy (anonymous):

@nar10

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

would would be the fraction of that? my answer choices are 5/7, 4 or 7/5, 4

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (anonymous):

woAH

OpenStudy (anonymous):

Use elimination. 2*(7x-2y)=2*(-3) 14x-4y=-6 Now 14x-4y=-6 -(14x+ y)=14 ------------- 0 -5y=-20 So y=4 For y=4 14x+y=14 14x+4=14 14x=10 x=10/14 x=5/7 x=5/7, y=4

OpenStudy (anonymous):

@nar10